[Brahms-l] Shift report 20050306 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun Mar 06 2005 - 08:11:29 EST
Shifter leader was SB
Shifters were 


1. Two hour beam continued from previous shift. Beam dumped at 2:20.
RHIC had a Cryogenic system problem since then which was recently solved and we just got a new store.
2. From previous shift
FS@8 deg @B2442, MRS@40 deg @A700
---------------------------------
13593  trig2(132k)  trig3(206k)
---------------------------------
>From today's night shift

13594  trig2(65k)  trig3(275k)
Non-TPC pedestal r13595
TPC pedestal r13596

3. No Detector problems

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Mar 6 08:11:45 2005

This archive was generated by hypermail 2.1.8 : Sun Mar 06 2005 - 08:11:49 EST